# Efficient stochastic simulation of gene regulatory networks using hybrid models of transcriptional bursting

### v1

@@ 1 Introduction
This paper is organized as follows. In Section 2 we introduce the model, and then detail the simulation algorithm in Section 3. The interest of this model is discussed in Section 4, and a proof of the algorithm is given in Section 5.
@@ 2 Mathematical model
[…]
where $G_{i}$, ${G_{i}}^{*}$, $M_{i}$ and $P_{i}$ denote inactive promoter (off), active promoter (on), mRNA and protein for gene $i$. These reactions describe the main two stages of gene expression, namely transcription (rate $s_{0,i}$) and translation (rate $s_{1,i}$), along with degradation of mRNA (rate $d_{0,i}$) and protein (rate $d_{1,i}$) molecules. Note that $G_{i}$ and ${G_{i}}^{*}$ satisfy a conservation relation and only one molecule of $G_{i}$ or ${G_{i}}^{*}$ is present at a time: for convenience we define $E_{i}\in\{0,1\}$ based on ${G_{i}}^{*}$. The particularity of the two-state model is that transcription can only occur when the promoter is active, i.e., $E_{i}=1$ (Fig. 1A). The bursty regime corresponds to $k_{\text{off},i}\gg k_{\text{on},i}$, meaning that promoter active periods are short compared to inactive periods (with $s_{0,i}\gg d_{0,i}$ such that $k_{\text{on},i}s_{0,i}/(k_{\text{off},i}d_{0,i})$ is fixed). In the limiting case, active periods become infinitely short so the promoter state is not needed anymore and the gene activity is characterized by its burst frequency $k_{\text{on},i}$ (Fig. 1B).
@@
$\left\{\begin{array}[]{rl}E_{i}(t)&:0\xrightarrow{k_{\text{on},i}}1,\,\;1\xrightarrow{k_{\text{off},i}}0\vspace{1.5mm}\\ {M_{i}}^{\prime}(t)&=s_{0,i}E_{i}(t)-d_{0,i}M_{i}(t)\vspace{1.5mm}\\ {P_{i}}^{\prime}(t)&=s_{1,i}M_{i}(t)-d_{1,i}P_{i}(t)\end{array}\right.$ (2)
for each gene $i\in\{1,\dots,n\}$, where we omit the dependence of $k_{\text{on},i}$ on $P$ for clarity. An example of stochastic trajectory is given in Fig. 2 (bottom left), to be compared with usual SSA (Fig. 2, top left). This is the model being considered in [34, 14, 4, 18]. It is also similar to that of [19], with two differences: (1) we describe explicitly mRNA levels (not only protein levels) and (2) we consider that interactions arise only from $k_{\text{on},i}$ and most importantly are not linear (i.e., $k_{\text{on},i}$ is a nonlinear, bounded function of $P$): this is an important aspect that is necessary to match real data (see Section 4).
Figure 2: Different mathematical frameworks for the stochastic gene expression model described in Fig. 1. The basic SSA (top left) corresponds to a continuous-time Markov chain implementing the stochastic mass action kinetics (1). The bursty version (top right) removes the promoter description and is driven by instantaneous bursts following a geometric distribution, while the hybrid version (bottom left) describes continuous levels of mRNA and protein, forming a PDMP. The model we consider here is another PDMP that combines these two simplifications (bottom right). The burst sizes then follow an exponential distribution as shown in Fig. 1B. To allow for comparison between models, the bursts are synchronized and the ranges of mRNA and protein in the four trajectories are identical (average $\langle M_{i}\rangle\approx 12$ copies/cell and $\langle P_{i}\rangle\approx 78$ copies/cell). Note that protein copy numbers are often much larger ($P_{i}\sim 10^{4}$ [41] or $10^{5}$ [1]) so that discrete trajectories become indistinguishable from their continuous counterparts.
@@
$\left\{\begin{array}[]{rl}M_{i}(t)&\xrightarrow{k_{\text{on},i}}M_{i}(t)+\mathcal{E}\Big{(}\frac{k_{\text{off},i}}{s_{0,i}}\Big{)}\vspace{0.5mm}\\ {M_{i}}^{\prime}(t)&=-d_{0,i}M_{i}(t)\vspace{1.5mm}\\ {P_{i}}^{\prime}(t)&=s_{1,i}M_{i}(t)-d_{1,i}P_{i}(t)\end{array}\right.$ (3)
for which an example of stochastic trajectory is given in Fig. 2 (bottom right), to be compared with bursty SSA (Fig. 2, top right). This model is used in [45, 28]. A similar model is considered in [31] (using a PDE approach) with the difference that here we also explicitly model mRNA in order to fit single-cell transcriptomic data (see also Section 4 about the interpretation of bursts). Analytical results have been obtained for a single gene with feedback [23]. Note that this is also precisely the limiting case considered in [34]. In absence of interaction between genes, the steady-state distribution of mRNA levels for each gene is a gamma distribution [23, 25], consistently with the derivation made in [34].
In the next section, we focus on exact simulation of model (3) as the simplest compromise for describing the transcriptional bursting property (Fig. 2). For mathematical completeness and with a view to proving the algorithm in Section 5, we mention that the model is characterized by its infinitesimal generator:
$\begin{array}[]{rll}Lf(y,z)&\displaystyle=\sum_{i=1}^{n}\left[-d_{0,i}y_{i}\frac{\partial}{\partial y_{i}}f(y,z)+(s_{1,i}y_{i}-d_{1,i}z_{i})\frac{\partial}{\partial z_{i}}f(y,z)\right.&\vspace{2mm}\\ &\qquad\displaystyle+\left.k_{\text{on},i}(z)\int_{0}^{+\infty}(f(y+he_{i},z)-f(y,z))b_{i}e^{-b_{i}h}\mathrm{d}h\right]\end{array}$ (4)
@@ 3 Simulation algorithm
{${}_{i}&=e^{-d_{0,i}t}y_{i}\\ [\varphi_{\text{P}}(y,z,t)]_{i}&=e^{-d_{1,i}t}z_{i}+\big{(}e^{-d_{1,i}t}-e^{-d_{0,i}t}\big{)}\frac{s_{1,i}}{d_{0,i}-d_{1,i}}y_{i}\end{aligned}\right.$
@@
and that will serve as a global “exponential clock” for computing jump times. From a computational perspective, the optimal value for lambda is the lowest possible one, which might not be available analytically. In practice, one can simply use $\lambda=\lambda_{1}+\cdots+\lambda_{n}$ where $k_{\text{on},i}(z)\leqslant\lambda_{i}$ are individual bounds. The interest of the simulation method also lies in the fact that the ODE system in (3) can be solved analytically.
@@
$p_{i}(z)=\begin{cases}\displaystyle 1-\frac{1}{\lambda}\sum_{i=1}^{n}k_{\text{on},i}(z)&\text{if $i=0$}\vspace{2mm}\\ \displaystyle\frac{k_{\text{on},i}(z)}{\lambda}&\text{if $1\leqslant i\leqslant n$}\vspace{1mm}\end{cases}$ (8)
@@
This is formalized in Algorithm 1 and a trajectory of the mRNA component $Y$ is represented in Fig. 3. We give here the algorithm to simulate one cell state $(Y(t),Z(t))$ but it can easily be modified to return cell trajectories collected along a time sequence $(t_{1},\dots,t_{k})$.
@@
2:$Y,Z\leftarrow Y_{0},Z_{0}$triangleright Initialize current state
3:$T\leftarrow 0$triangleright Initialize current jump time
@@
7:     $U\leftarrow\mathcal{E}(\lambda)$triangleright Draw waiting time using (7)
8:     $Y,Z\leftarrow\varphi(Y_{\text{old}},Z_{\text{old}},U)$triangleright Apply the deterministic flow (5)
9:     $i\leftarrow\mathcal{P}(Z)$triangleright Draw $i$ with probability given by (8)
@@
11:         $Y[i]\leftarrow Y[i]+\mathcal{E}(b_{i})$triangleright Apply jump to mRNA level of gene $i$
@@
13:     $T\leftarrow T+U$triangleright Update current jump time
@@
15:return $\varphi(Y_{\text{old}},Z_{\text{old}},t-T_{\text{old}})$triangleright Extend to final time
@@
In some circumstances, computational cost can still be improved. For example, we used the bound $\lambda$ over $\sum_{i}k_{\text{on},i}$, but a bound over $\sum_{i}k_{\text{on},i}\circ\varphi_{\text{P}}$ may be sharper, making it possible to reduce the number of phantom jumps. Note that the flow (5) has a unique stable attractor: combined with the fact that the jump rate is continuous, one can adapt the algorithm to unbounded jump rates. This requires computing an optimized bound at each iteration, as made for example in the implementation of the Harissa package [13]. Furthermore, we made a distinction here between $Y(t)$, which is impacted by jumps, and $Z(t)$, which is involved in interactions, but Algorithm 1 can be directly adapted to cases with more variables and other role combinations.
Figure 3: Illustration of Algorithm 1 in the case $n=2$, showing the iterative construction for three jump times: here the first jump applies to gene $1$, the second jump is rejected, and the third jump applies to gene $2$. For simplicity, only the mRNA component is shown; the protein component is continuous and derived directly from the corresponding deterministic flow (i.e., using the analytical solution of third equation in (3)).
@@ 4 Discussion
Contrary to the usual SSA, which becomes costly when simulating single-cell gene expression data [5], the PDMP model presented here is much faster to simulate (see, e.g., a factor of 200 in [14]). Mathematically, it can be justified as a rigorous hybrid limit of the discrete formalism when mRNA and protein levels span an important range [8, 9, 6], consistently with biological observations (see Fig. 2 and [41, 42, 1]). We also mention another interesting link: using the formalism of Poisson representation [11], it is possible to show that in absence of interactions, the discrete and hybrid models are in fact related by an exact intertwining relation, with no need for considering a limit regime. Namely, the distribution from the SSA model can then be written as a mixture of Poisson distributions, whose time-dependent mixing distribution is that of the PDMP counterpart [16, 12]. This type of result is of major interest when distinguishing measurement and expression models for single-cell data analysis [39].
@@
As an illustration, we show in Fig. 4 an example of basic toggle switch network, which generates a bistable pattern emerging from interactions. We believe this example is interesting because a new “two-state” layer appears, which seems to be consistent with previous diffusion modeling [15] while the authors were interpreting the two states as promoter states, possibly from a confusion of time scales and the fact that they had no single-cell data. The interpretation here is that bursts are transcriptional, with proteins tending to buffer these bursts, making the pattern robust.
@@ 5 Proof
In this section we present an explicit construction of the process corresponding to Algorithm 1 and illustrated in Fig. 3. We then prove that the infinitesimal generator of the resulting process is that of the PDMP model (4). This is sufficient to conclude that obtained trajectories are exact. The reader can refer to [2] for additional details on the underling theory.
@@
Algorithm 1 leads to the following construction. We first define a discrete-time Markov chain $(Y_{k},Z_{k})_{k\geqslant 0}\in\mathbb{R}_{+}^{n}\times\mathbb{R}_{+}^{n}$ which corresponds to the right limit of the process $(Y(t),Z(t))_{t\geqslant 0}\in\mathbb{R}_{+}^{n}\times\mathbb{R}_{+}^{n}$ at specific (random) time points. These time points are given by a homogeneous Poisson process $(N(t))_{t\geqslant 0}$ with intensity $\lambda$. Let $(T_{k})_{k\geqslant 0}$ denote its jump time sequence and $(U_{k})_{k\geqslant 1}$ its inter-arrival times. Let $Y_{0}$ and $Z_{0}$ be random variables independent of $(N(t))_{t\geqslant 0}$. For $k\geqslant 0$, we recursively define
{$Y_{k}&=\varphi_{\text{M}}(Y_{k-1},Z_{k-1},U_{k})+S_{k}\\ Z_{k}&=\varphi_{\text{P}}(Y_{k-1},Z_{k-1},U_{k})\end{aligned}\right.$
where $S_{k}$ is a random vector constructed as follows: at each step $k$, we draw a random index $I\in\{0,1,\dots,n\}$ according to the probability distribution $\mathcal{P}(Z_{k})$ defined at (8). When $I\neq 0$, we consider an additional variable $E$ from exponential distrobution $\mathcal{E}(b_{I})$. We set
$S_{k}=\begin{cases}0&\text{if $I=0$}\vspace{2mm}\\ Ee_{i}&\text{if $1\leqslant i\leqslant n$}\vspace{1mm}\end{cases}$

### v2

@@ 1 Introduction
This paper is organized as follows. In Section˜2 we introduce the model, and then detail the simulation algorithm in Section˜3. The interest of this model is discussed in Section˜4, and a proof of the algorithm is given in Section˜5.
@@ 2 Mathematical model
$\begin{array}[]{ccc}G_{i}\xrightleftharpoons[k_{\text{off},i}]{k_{\text{on},i}}{G_{i}}^{*},&{G_{i}}^{*}\xrightarrow[]{s_{0,i}}{G_{i}}^{*}+M_{i},&M_{i}\xrightarrow[]{d_{0,i}}\varnothing,\\ &M_{i}\xrightarrow[]{s_{1,i}}M_{i}+P_{i},&P_{i}\xrightarrow[]{d_{1,i}}\varnothing,\end{array}$ (1)
where $G_{i}$, $G_{i}$$^{*}$, $M_{i}$ and $P_{i}$ denote inactive promoter (off), active promoter (on), mRNA and protein for gene $i$. These reactions describe the main two stages of gene expression, namely transcription (rate $s_{0,i}$) and translation (rate $s_{1,i}$), along with degradation of mRNA (rate $d_{0,i}$) and protein (rate $d_{1,i}$) molecules. Note that $G_{i}$ and $G_{i}$$^{*}$ satisfy a conservation relation and only one molecule of $G_{i}$ or $G_{i}$$^{*}$ is present at a time: for convenience we define $E_{i}\in\{0,1\}$ based on $G_{i}$$^{*}$. The particularity of the two-state model is that transcription can only occur when the promoter is active, i.e., $E_{i}=1$ (Fig.˜1A). The bursty regime corresponds to $k_{\text{off},i}\gg k_{\text{on},i}$, meaning that promoter active periods are short compared to inactive periods (with $s_{0,i}\gg d_{0,i}$ such that $k_{\text{on},i}s_{0,i}/(k_{\text{off},i}d_{0,i})$ is fixed). In the limiting case, active periods become infinitely short so the promoter state is not needed anymore and the gene activity is characterized by its burst frequency $k_{\text{on},i}$ (Fig.˜1B).
@@
$\left\{\begin{array}[]{rl}E_{i}(t)&\hskip-5.69054pt:0\xrightarrow{k_{\text{on},i}}1,\,\;1\xrightarrow{k_{\text{off},i}}0\vskip 4.2679pt\\ {M_{i}}^{\prime}(t)&\hskip-5.69054pt=s_{0,i}E_{i}(t)-d_{0,i}M_{i}(t)\vskip 4.2679pt\\ {P_{i}}^{\prime}(t)&\hskip-5.69054pt=s_{1,i}M_{i}(t)-d_{1,i}P_{i}(t)\end{array}\right$. (2)
for each gene $i\in\{1,\dots,n\}$, where we omit the dependence of $k_{\text{on},i}$ on $P$ for clarity. An example of stochastic trajectory is given in Fig.˜2 (bottom left), to be compared with usual SSA (Fig.˜2, top left). This is the model being considered in [34, 14, 4, 18]. It is also similar to that of [19], with two differences: (1) we describe explicitly mRNA levels (not only protein levels) and (2) we consider that interactions arise only from $k_{\text{on},i}$ and most importantly are not linear (i.e., $k_{\text{on},i}$ is a nonlinear, bounded function of $P$): this is an important aspect that is necessary to match real data (see Section˜4).
Figure 2: Different mathematical frameworks for the stochastic gene expression model described in Fig.˜1. The basic SSA (top left) corresponds to a continuous-time Markov chain implementing the stochastic mass action kinetics (1). The bursty version (top right) removes the promoter description and is driven by instantaneous bursts following a geometric distribution, while the hybrid version (bottom left) describes continuous levels of mRNA and protein, forming a PDMP. The model we consider here is another PDMP that combines these two simplifications (bottom right). The burst sizes then follow an exponential distribution as shown in Fig.˜1B. To allow for comparison between models, the bursts are synchronized and the ranges of mRNA and protein in the four trajectories are identical (average $\langle M_{i}\rangle\approx 12$ copies/cell and $\langle P_{i}\rangle\approx 78$ copies/cell). Note that protein copy numbers are often much larger ($P_{i}\sim 10^{4}$ [41] or $10^{5}$ [1]) so that discrete trajectories become indistinguishable from their continuous counterparts.
@@
$\left\{\begin{array}[]{rl}M_{i}(t)&\hskip-5.69054pt\xrightarrow{k_{\text{on},i}}M_{i}(t)+\mathcal{E}\Big{(}\frac{k_{\text{off},i}}{s_{0,i}}\Big{)}\vskip 1.42262pt\\ {M_{i}}^{\prime}(t)&\hskip-5.69054pt=-d_{0,i}M_{i}(t)\vskip 4.2679pt\\ {P_{i}}^{\prime}(t)&\hskip-5.69054pt=s_{1,i}M_{i}(t)-d_{1,i}P_{i}(t)\end{array}\right$. (3)
for which an example of stochastic trajectory is given in Fig.˜2 (bottom right), to be compared with bursty SSA (Fig.˜2, top right). This model is used in [45, 28]. A similar model is considered in [31] (using a PDE approach) with the difference that here we also explicitly model mRNA in order to fit single-cell transcriptomic data (see also Section˜4 about the interpretation of bursts). Analytical results have been obtained for a single gene with feedback [23]. Note that this is also precisely the limiting case considered in [34]. In absence of interaction between genes, the steady-state distribution of mRNA levels for each gene is a gamma distribution [23, 25], consistently with the derivation made in [34].
In the next section, we focus on exact simulation of model (3) as the simplest compromise for describing the transcriptional bursting property (Fig.˜2). For mathematical completeness and with a view to proving the algorithm in Section˜5, we mention that the model is characterized by its infinitesimal generator:
$\hskip-5.69054pt\begin{array}[]{rll}Lf(y,z)&\displaystyle\hskip-5.69054pt=\sum_{i=1}^{n}\left[-d_{0,i}y_{i}\frac{\partial}{\partial y_{i}}f(y,z)+(s_{1,i}y_{i}-d_{1,i}z_{i})\frac{\partial}{\partial z_{i}}f(y,z)\right.&\hskip-8.53581pt\vskip 5.69054pt\\ &\qquad\displaystyle+\left.k_{\text{on},i}(z)\int_{0}^{+\infty}(f(y+he_{i},z)-f(y,z))b_{i}e^{-b_{i}h}\mathrm{d}h\right]\end{array}$ (4)
@@ 3 Simulation algorithm
$\left\{\begin{aligned} {}_{i}&=e^{-d_{0,i}t}y_{i}\\ [\varphi_{\text{P}}(y,z,t)]_{i}&=e^{-d_{1,i}t}z_{i}+\big{(}e^{-d_{1,i}t}-e^{-d_{0,i}t}\big{)}\frac{s_{1,i}}{d_{0,i}-d_{1,i}}y_{i}\end{aligned}\right$.
@@
and that will serve as a global “exponential clock” for computing jump times. From a computational perspective, the optimal value for $\lambda$ is the lowest possible one, which might not be available analytically. In practice, one can simply use $\lambda=\lambda_{1}+\cdots+\lambda_{n}$ where $k_{\text{on},i}(z)\leqslant\lambda_{i}$ are individual bounds. The interest of the simulation method also lies in the fact that the ODE system in (3) can be solved analytically.
@@
$p_{i}(z)=\begin{cases}\displaystyle 1-\frac{1}{\lambda}\sum_{i=1}^{n}k_{\text{on},i}(z)&\text{if $i=0$}\vskip 5.69054pt\\ \displaystyle\frac{k_{\text{on},i}(z)}{\lambda}&\text{if $1\leqslant i\leqslant n$}\vskip 2.84526pt\end{cases}$ (8)
@@
This is formalized in Algorithm˜1 and a trajectory of the mRNA component $Y$ is represented in Fig.˜3. We give here the algorithm to simulate one cell state $(Y(t),Z(t))$ but it can easily be modified to return cell trajectories collected along a time sequence $(t_{1},\dots,t_{k})$.
@@
[…]
@@
7:  $U\leftarrow\mathcal{E}(\lambda)$ $\triangleright$ Draw waiting time using (7)
8:  $Y,Z\leftarrow\varphi(Y_{\text{old}},Z_{\text{old}},U)$ $\triangleright$ Apply the deterministic flow (5)
9:  $i\leftarrow\mathcal{P}(Z)$ $\triangleright$ Draw $i$ with probability given by (8)
@@
11:   $Y[i]\leftarrow Y[i]+\mathcal{E}(b_{i})$ $\triangleright$ Apply jump to mRNA level of gene $i$
@@
13:  $T\leftarrow T+U$ $\triangleright$ Update current jump time
@@
15:return $\varphi(Y_{\text{old}},Z_{\text{old}},t-T_{\text{old}})$ $\triangleright$ Extend to final time
@@
In some circumstances, computational cost can still be improved. For example, we used the bound $\lambda$ over $\sum_{i}k_{\text{on},i}$, but a bound over $\sum_{i}k_{\text{on},i}\circ\varphi_{\text{P}}$ may be sharper, making it possible to reduce the number of phantom jumps. Note that the flow (5) has a unique stable attractor: combined with the fact that the jump rate is continuous, one can adapt the algorithm to unbounded jump rates. This requires computing an optimized bound at each iteration, as made for example in the implementation of the Harissa package [13]. Furthermore, we made a distinction here between $Y(t)$, which is impacted by jumps, and $Z(t)$, which is involved in interactions, but Algorithm˜1 can be directly adapted to cases with more variables and other role combinations.
Figure 3: Illustration of Algorithm˜1 in the case $n=2$, showing the iterative construction for three jump times: here the first jump applies to gene $1$, the second jump is rejected, and the third jump applies to gene $2$. For simplicity, only the mRNA component is shown; the protein component is continuous and derived directly from the corresponding deterministic flow (i.e., using the analytical solution of third equation in (3)).
@@ 4 Discussion
Contrary to the usual SSA, which becomes costly when simulating single-cell gene expression data [5], the PDMP model presented here is much faster to simulate (see, e.g., a factor of 200 in [14]). Mathematically, it can be justified as a rigorous hybrid limit of the discrete formalism when mRNA and protein levels span an important range [8, 9, 6], consistently with biological observations (see Fig.˜2 and [41, 42, 1]). We also mention another interesting link: using the formalism of Poisson representation [11], it is possible to show that in absence of interactions, the discrete and hybrid models are in fact related by an exact intertwining relation, with no need for considering a limit regime. Namely, the distribution from the SSA model can then be written as a mixture of Poisson distributions, whose time-dependent mixing distribution is that of the PDMP counterpart [16, 12]. This type of result is of major interest when distinguishing measurement and expression models for single-cell data analysis [39].
@@
As an illustration, we show in Fig.˜4 an example of basic toggle switch network, which generates a bistable pattern emerging from interactions. We believe this example is interesting because a new “two-state” layer appears, which seems to be consistent with previous diffusion modeling [15] while the authors were interpreting the two states as promoter states, possibly from a confusion of time scales and the fact that they had no single-cell data. The interpretation here is that bursts are transcriptional, with proteins tending to buffer these bursts, making the pattern robust.
@@ 5 Proof
In this section we present an explicit construction of the process corresponding to Algorithm˜1 and illustrated in Fig.˜3. We then prove that the infinitesimal generator of the resulting process is that of the PDMP model (4). This is sufficient to conclude that obtained trajectories are exact. The reader can refer to [2] for additional details on the underling theory.
@@
Algorithm˜1 leads to the following construction. We first define a discrete-time Markov chain $(Y_{k},Z_{k})_{k\geqslant 0}$ where $(Y_{k},Z_{k})\in\mathbb{R}_{+}^{n}\times\mathbb{R}_{+}^{n}$, which corresponds to the right limit of the process $(Y(t),Z(t))_{t\geqslant 0}$ at specific (random) time points. These time points are given by a homogeneous Poisson process $(N(t))_{t\geqslant 0}$ with intensity $\lambda$. Let $(T_{k})_{k\geqslant 0}$ denote its jump time sequence and $(U_{k})_{k\geqslant 1}$ its inter-arrival times. Let $Y_{0}$ and $Z_{0}$ be random variables independent of $(N(t))_{t\geqslant 0}$. For $k\geqslant 0$, we recursively define
$\left\{\begin{aligned} Y_{k}&=\varphi_{\text{M}}(Y_{k-1},Z_{k-1},U_{k})+S_{k}\\ Z_{k}&=\varphi_{\text{P}}(Y_{k-1},Z_{k-1},U_{k})\end{aligned}\right$.
where $S_{k}$ is a random vector constructed as follows: at each step $k$, we draw a random index $i\in\{0,1,\dots,n\}$ according to the probability distribution $\mathcal{P}(Z_{k})$ defined at (8). When $i\neq 0$, we consider an additional variable $E$ with exponential distribution $\mathcal{E}(b_{i})$. We set
$S_{k}=\begin{cases}0&\text{if $i=0$}\vskip 5.69054pt\\ Ee_{i}&\text{if $1\leqslant i\leqslant n$}\vskip 2.84526pt\end{cases}$